LABORATORY DETECTION OF IZnCH_{3} (X^{1}A_{1}) : FURTHER EVIDENCE FOR ZINC INSERTION

\underline{MATTHEW P. BUCCHINO}, Department of Chemistry and Biochemistry, University of Arizona, Tucson, Arizona 85721; JUSTIN P. YOUNG, Department of Chemistry, Canisius College, Buffalo, New York 14208; PHIL M. SHERIDAN, Department of Chemistry, Canisius College, Buffalo, New York 14208; and LUCY M. ZIURYS, Department of Chemistry and Biochemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, Arizona 85721.

Millimeter-wave direct absorption techniques were used to record the pure rotational spectrum of IZnCH_{3} (X^{1}A_{1}). This species was produced by the reaction of zinc vapor with ICH_{3} in the presence of a DC discharge. Rotational transitions ranging from J = 109 → 108 to J = 122 → 121 were recorded for I^{64}ZnCH_{3} and I^{66}ZnCH_{3} in the frequency range of 250−290 GHz. The Ka = 0−4 components were measured for each transition, with the K-ladder structure and nuclear spin statistics indicative of a symmetric top. As with HZnCH_{3} (X^{1}A_{1}), the detection of IZnCH_{3} provides further evidence for a zinc insertion process.